OpenStudy (anonymous):

Find the measures of two angles, one positive and one negative, that are coterminal with pi/5.

OpenStudy (anonymous):

So what are coterminal angles? Two angles are coterminal if they have the same terminal side. In geometry or picture view, it is easy to see a coterminal angle. But from mathematical point of view, we can see it as following: A is an angle Then coterminal of A should be: \[A \pm k.2\pi\] with k is an integter number (k can be any negative like -1 or positive numbers +1, +2, ....) In your question, \[A = \pi/5 \rightarrow B = A \pm k.2\pi = \pi/5 \pm k.2\pi\] For example, if you want a positive angle, then you may try k = +1, \[B = 11\pi/5\] A negative angle, you may try k = -1, \[B = -9\pi/5\]

OpenStudy (anonymous):

I have answer choices: \[ 11\pi/5;-9\pi/5, \pi/5 + 360degrees; \pi/5 - 360degrees, 6\pi/5;-4\pi/5, 11\pi/5;-\pi/5\]

OpenStudy (anonymous):

One of my choices was 11pi/5, and -9/5. Is that right?

OpenStudy (anonymous):

yes, you right, :) there are many options, but the easiest one is to assign k = +/- 1

OpenStudy (anonymous):

pi/5 + 2pi = pi/5 + 10pi/5 = 11pi/5 and pi/5 - 2pi

OpenStudy (anonymous):

Thank you @DSS , but I'm confused on how pi/5 + 2pi = 11pi/5. Can you explain that?

OpenStudy (anonymous):

\[\frac{ 1 }{ 5 } + 2 = \frac{ 1 + 10 }{ 5 }\]

OpenStudy (anonymous):

Is that because you faound a common denominator, so you really did 1/5 + 10/5? Because it started as 2/1, and you multiplied by 5?

OpenStudy (anonymous):

Yes, actually, this is very basic pre-algebra

OpenStudy (anonymous):

Okay, then I understand it now. That website she linked made it simple. I don't know why my class website can't just explain it that way. They try to make it so confusing..

OpenStudy (anonymous):

Thank you guys again. I only have two more questions! haha. Then maybe I will be able to do all of this on my own. :)